# A Non-Abelian Approach to Riemann Surfaces

### 04 Mar 2026 v1

@@ 1. Introduction
Following Klein’s account in [16, Chapter III, §6], this invariant goes back to Lagrange (see [17]). Since the Schwarzian derivative is invariant under Möbius transformations, the quantity $S(\tau)$ is well defined, independent of the chosen basis. Dedekind’s computation asserts that $S(\tau)$ is in fact the rational function
@@
We discuss the following three applications of the non-abelian Schwarzian formalism.
@@
    Periods of Riemann surfaces of higher genus (Sections 12 and 13). Periods in a holomorphic family satisfy Picard–Fuchs equations arising from the Gauss–Manin connection (see [15, 14, 9]). For a one-parameter family of genus $g$ curves, the classical reduction typically yields a non-canonical scalar equation of order $2g$ [14]. For a generic family (Definition 12.1) we instead construct canonically a second-order equation with $g\times g$ matrix coefficients on the Hodge bundle, whose solutions are period vectors; the associated matrix Schwarzian produces characteristic invariants generalizing Dedekind’s genus-one computation.
@@
    Periods of higher-dimensional varieties: cubic threefolds (Section 14). For a one-parameter family of smooth cubic threefolds in $\mathbb{P}^{4}$, the Gauss–Manin connection gives a first-order system on $H^{3}$ of rank $10$. Our approach produces a canonical second-order system on the rank-55 Hodge bundle $E=H^{2,1}$, yielding Schwarzian-type invariants for this higher-dimensional variation of Hodge structure.
@@
    Mass–spring systems (Section 15). Mass–spring systems naturally give rise to second-order ordinary differential equations with matrix-valued coefficients. This example clarifies the philosophy of the paper: time is treated as a curve and the coefficients become geometric data along that curve, transforming coherently under changes of local time coordinate.
@@
Part IV (Sections 12–15) treats periods of Riemann surfaces and cubic threefolds from a Schwarzian perspective.
Part V (Sections 16–19) discusses the classical and quantum damped mass–spring system in mechanics and its relation to the Schwarzian.
@@ 2. Second order scalar ODE
The following classical computation is standard (see, for example, [11, § IV.2]). For a classical discussion of the projective viewpoint on second-order equations, see also Klein’s Lectures on the Icosahedron [16, Chapter III, §7].
@@ 4. Classical differentials and connections
Recall that $K_{X}$ denotes the canonical line bundle of holomorphic 11-forms on $X$. Fix an integer $m$. A meromorphic connection on $K_{X}^{m}$ may be encoded locally by meromorphic functions $a_{i}\in\mathcal{M}(U_{i})$ satisfying on overlaps
@@ 5. Quantum differentials and connections
Here II denotes the identity operator on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},V)$, scalar functions act by pointwise multiplication, and the transported operator is
@@
on the open subset of $U_{i}$ where $\Psi_{i}(f)$ is pointwise invertible. Then $A=\{A_{i}\}$ is a quantum connection of eccentricity 11 with values in $\operatorname{End}(V)$.
@@
one obtains a scalar quantum connection of eccentricity 11.
@@ 6. Curvature
Let AA be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. For $m\in\mathbb{Z}$ and a quantum $m$-differential $\Psi$ with values in $V$, define the quantum covariant derivative $\nabla_{A}\Psi$ locally by
@@
If AA has eccentricity $e\neq 0$ and $\Psi$ is a quantum $m$-differential, then $\nabla_{A}\Psi$ is a quantum $(m+1)$-differential.
@@
Even if AA and $\Psi$ are classical in the sense of the previous example, the operator $\nabla_{A}\Psi$ need not be classical: the derivative term $\Psi_{i}(f)^{\prime}$ typically produces an $f^{\prime}$-term. In the classical setting one instead defines the coefficient-level covariant derivative by
@@
We briefly recall the classical Cartan structural equation for a connection with values in a Lie algebra. Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A $\mathfrak{g}$-valued connection 11-form on a manifold $M$ is a $\mathfrak{g}$-valued differential 11-form $A\in\Omega^{1}(M,\mathfrak{g})$. Its curvature is the $\mathfrak{g}$-valued 22-form
@@
Let AA be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. The quantum curvature of AA is the family of local operators
@@
Let AA be a quantum connection of eccentricity $e\neq 0$ with values in $\operatorname{End}(V)$. On an overlap $U_{i}\cap U_{j}$ with change of coordinate $\lambda=\lambda_{ij}$, one has
@@
where $S(\lambda)$ is the Schwarzian derivative of $\lambda$, and II is the identity endomorphism.
@@
Consider the scalar quantum 11-differential $\Psi(f)=f^{\prime}$. On the open set where $f^{\prime}\neq 0$, the Maurer–Cartan construction yields the scalar quantum connection $A(f)=f^{\prime\prime}/f^{\prime}$. Its quantum curvature is
@@
Let AA be a quantum connection with values in $\operatorname{End}(V)$ and eccentricity $e$. For $g\in G$ define
@@
We recall that the appearance of an affine “connection term” in the transformation law is classical; see Klein [16, Chapter III, §7] for the corresponding discussion in the genus one setting.
@@
If AA is a quantum connection of eccentricity $e$, then $g\bullet A$ is again a quantum connection of eccentricity $e$. Moreover, $(g,A)\mapsto g\bullet A$ defines an action of $G$.
@@
Here II is the identity operator, the matrices $g_{ij}$ act by pointwise multiplication, and primes denote differentiation with respect to the local coordinate on $U_{j}$.
@@
The anomaly term $eS(\lambda)I$ is scalar (proportional to II). Since elements of $G$ act by conjugation and scalar endomorphisms commute with all elements of $\operatorname{End}(V)$, the relation $F_{A_{2}}=gF_{A_{1}}g^{-1}$ is independent of the choice of local coordinates. Thus the definition is well defined.
@@ 7. Second order quantum ODE
whose coefficients depend on the chosen coordinate. In the quantum setting we take the solution to be a quantum 0-differential $\Psi$, and we allow the coefficients to be quantum tensors (connections and 22-differentials) so that the notion becomes intrinsic on $X$.
@@
as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j})$. In particular, $q$ is a quantum 22-differential, while $p$ transforms as a quantum connection of eccentricity 11.
@@
By Lemma 7.1, the coefficient $p$ transforms as a quantum connection of eccentricity 11. Hence we may view $p$ itself as a connection and form the corresponding covariant derivative $\nabla_{p}$ acting on quantum differentials.
@@
be the scalar quantum connection (of eccentricity 11) on $U_{i}$. Let
@@
be the quantum 22-differential given by squaring the first derivative. Fix a constant $c\in\mathbb{C}$, and set
@@
where $\Psi$ is a quantum 0-differential and $q$ is a quantum 22-differential. Equivalently, since for $e=\tfrac{1}{2}$ and $m=0$ one has
@@ 7.1. Main Lemma
We are now ready to state our main lemma. Suppose we have an invertible solution $g$ of this equation with values in $\operatorname{End}(V)$, then the corresponding gauge transform of AA has curvature $F_{A}-q$ up to conjugation.
@@
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $g$ be an $\operatorname{End}(V)$-valued quantum 0-differential such that for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$ the meromorphic function $g_{i}(f)\in\mathcal{M}(U_{i},\operatorname{End}(V))$ satisfies $\det(g_{i}(f))\not\equiv 0$ on $U_{i}$. Assume that for every $i$ and every $f\in\mathcal{M}(U_{i})$ one has, on the open subset where $g_{i}(f)$ is pointwise invertible,
@@
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. For the quantum second-order equation
@@
while $q$ transforms tensorially as a quantum 22-differential. Hence
@@
In particular, $\mathcal{S}_{A,q}$ has the same anomaly under coordinate change as the curvature (up to the normalizing factor 22).
@@ 8. The gauge and $\star$ actions
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. The second-order equation (7.0) depends on the choice of the pair $(A,q)$, while its quantum Schwarzian $\mathcal{S}_{A,q}=2(F_{A}-q)$ is a more intrinsic quantity. We now describe a natural action of quantum 11-differentials on such pairs which preserves $F_{A}-q$ (and hence $\mathcal{S}_{A,q}$).
@@
Let $u$ be an $\operatorname{End}(V)$-valued quantum 11-differential on $X$. We define
@@
The operation $\star$ defines an action of the additive group of $\operatorname{End}(V)$-valued quantum 11-differentials on the set of pairs $(A,q)$ with $e$. Equivalently, for quantum 11-differentials $u$ and $v$ one has
@@
This is a direct computation using bilinearity of the products and linearity of the derivative. Writing $(A^{\prime},q^{\prime}):=u\star(A,q)$, we have $A^{\prime}=A+u$ and $q^{\prime}=q-uA-Au-u^{2}+u^{\prime}$. Hence
$v\star(A^{\prime},q^{\prime})$ $=\bigl(A+u+v,\;q-uA-Au-u^{2}+u^{\prime}-v(A+u)-(A+u)v-v^{2}+v^{\prime}\bigr)$
@@ 8.1. Main Theorem
Let $g\in G$ be a gauge transformation (Definition 6.7). Let AA be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Define
@@
Then $A^{g}$ is again a quantum connection of eccentricity $e=\tfrac{1}{2}$, and $q^{g}$ is an $\operatorname{End}(V)$-valued quantum 22-differential. Moreover, if $\Psi$ is a (quantum) 0-differential solving
@@
and subtracting $q^{g}$ yields $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$. Multiplying by 22 gives the stated transformation of $\mathcal{S}_{A,q}$. Conjugation invariance of characteristic polynomial coefficients and traces is immediate.
qedhere ∎
@@
It is important that the transformation of $q$ in Lemma 8.4 is not given by tensorial conjugation alone. In general the curvature $F_{A}=A^{\prime}-A^{2}$ is not gauge covariant under the action $A\mapsto g\bullet A=gAg^{-1}+g^{\prime}g^{-1}$ when AA is $\operatorname{End}(V)$-valued and noncommutative. The additional terms $2g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime}$ in $q^{g}$ are exactly what is needed so that the Schwarzian combination $F_{A}-q$ transforms by conjugation, and hence $\mathcal{S}_{A,q}=2(F_{A}-q)$ is the natural gauge-theoretic invariant of the second order equation.
@@
Let AA be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $u$ be an $\operatorname{End}(V)$-valued quantum 11-differential and let $g\in G$ be a gauge transformation. Write
$u\star(A,q)=(A+u,\;\widetilde{q})\qquad\text{with}\qquad\widetilde{q}:=q-uA-Au-u^{2}+u^{\prime}.$
[…]
$\bigl((A+u)^{g},\widetilde{q}^{g}\bigr)\;=\;u^{g}\star(A^{g},q^{g}).$
Equivalently, the gauge group acts by automorphisms of the $\star$-action, so $\star$-orbits of pairs $(A,q)$ are preserved under gauge transformations.
@@
The first components agree:
$(A+u)^{g}=g^{-1}(A+u)g-g^{-1}g^{\prime}=(g^{-1}Ag-g^{-1}g^{\prime})+g^{-1}ug=A^{g}+u^{g}.$
For the second component, it is enough to compare the gauge-covariant combination $F_{A}-q$. By Lemma 8.3 one has $F_{A+u}-\widetilde{q}=F_{A}-q$, and by Lemma 8.4 this implies
$F_{(A+u)^{g}}-\widetilde{q}^{g}=g^{-1}(F_{A}-q)g.$
On the other hand, Lemma 8.4 gives $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$, and applying Lemma 8.3 to the pair $(A^{g},q^{g})$ and the 11-differential $u^{g}$ yields
$F_{A^{g}+u^{g}}-q_{u^{g}\star(A^{g},q^{g})}=F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g.$
Since $(A+u)^{g}=A^{g}+u^{g}$, the two pairs $\bigl((A+u)^{g},\widetilde{q}^{g}\bigr)$ and $u^{g}\star(A^{g},q^{g})$ have the same first component and the same value of $F_{A}-q$; therefore their second components coincide, proving the claim. ∎
@@
The action $\star$ is transitive on the choice of the connection component: if $A_{1}$ and $A_{2}$ are two eccentricity $e=\tfrac{1}{2}$ quantum connections, then $u:=A_{2}-A_{1}$ is a quantum 11-differential and $A_{2}=A_{1}+u$. Thus $\star$ parametrizes different presentations $(A,q)$ of the same quantum Schwarzian. In the classical scalar case this recovers the familiar fact that a projective connection can be represented by an affine connection modulo addition of a 11-form.
Example 8.8 (The quantum Schwarzian in Example 4.3).
Consider the scalar equation of Example 4.3,
$Y^{\prime\prime}=\frac{f^{\prime\prime}}{f^{\prime}}\,Y^{\prime}+c^{2}(f^{\prime})^{2}\,Y,$
and rewrite it in the eccentricity $e=\tfrac{1}{2}$ form
$Y^{\prime\prime}=2p\,Y^{\prime}+q\,Y$
by setting
$p=\frac{1}{2}\,\frac{f^{\prime\prime}}{f^{\prime}},\qquad q=c^{2}(f^{\prime})^{2}.$
Then the corresponding (scalar) quantum Schwarzian is
$\mathcal{S}_{p,q}=2\,(F_{p}-q)=2\bigl(p^{\prime}-p^{2}-q\bigr)=\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}-2c^{2}(f^{\prime})^{2}=S(f)-2c^{2}(f^{\prime})^{2}.$
On the other hand, the two linearly independent local solutions found in Example 4.3 are $y_{1}=e^{cf}$ and $y_{2}=e^{-cf}$, hence
$w:=\frac{y_{1}}{y_{2}}=e^{2cf}.$
We now compute $S(w)$ in two different ways and compare it with the quantum Schwarzian.
Method 1: from the differential equation. By Proposition 2.1, the ratio $w=y_{1}/y_{2}$ satisfies
$S(w)=2\bigl(p^{\prime}-p^{2}-q\bigr)=\mathcal{S}_{p,q}.$
Method 2: by the Schwarzian chain rule. Using the Schwarzian chain rule $S(h\circ f)=(S(h)\circ f)(f^{\prime})^{2}+S(f)$ and the direct computation $S(e^{2ct})=-2c^{2}$, we obtain
$S(w)=S(e^{2cf})=S(f)-2c^{2}(f^{\prime})^{2}.$
This agrees with the expression for $\mathcal{S}_{p,q}$ above, and thus recovers the identity $S(y_{1}/y_{2})=\mathcal{S}_{p,q}$ in this concrete example.
@@
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Suppose $g$ is an $\operatorname{End}(V)$-valued quantum 0-differential such that $\det(g_{i}(f))\not\equiv 0$ for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$, and assume that $g$ satisfies the second-order equation
@@
Let AA be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum 22-differential. Let $g_{1}$ and $g_{2}$ be two invertible (classical) $\mathrm{GL}(V)$-valued solutions of

### 14 Mar 2026 v2

@@ 1. Introduction
Following Klein’s account in [18, Chapter III, §6], this invariant goes back to Lagrange (see [19]). Since the Schwarzian derivative is invariant under Möbius transformations, the quantity $S(\tau)$ is well defined, independent of the chosen basis. Dedekind’s computation asserts that $S(\tau)$ is in fact the rational function
@@
Some of the ideas developed here continue in a higher-order direction in recent work of Jafari [15], where the reparametrization formalism is extended to higher Schwarzian cocycles and noncommutative Wilczynski-type invariants.
@@
We discuss three applications of the non-abelian Schwarzian formalism. The same formalism produces canonical second-order systems in three different settings: periods of curves, periods of higher-dimensional varieties, and mechanical systems. The higher-genus case is the main geometric application: it replaces a non-canonical scalar Picard–Fuchs equation of order $2g$ by a canonical second-order equation with $g\times g$ matrix coefficients on the Hodge bundle.
@@
    Periods of Riemann surfaces of higher genus (Sections 12 and 13). Periods in a holomorphic family satisfy Picard–Fuchs equations arising from the Gauss–Manin connection (see [17, 16, 10]). For a one-parameter family of genus $g$ curves, the classical reduction typically yields a non-canonical scalar equation of order $2g$ [16]; for a generic family (Definition 12.1) we instead construct canonically a second-order equation with $g\times g$ matrix coefficients on the Hodge bundle, whose solutions are period vectors. This canonical system is the natural higher-genus analogue of Dedekind’s genus-one picture, and its associated matrix Schwarzian produces the corresponding characteristic invariants.
@@
    Periods of higher-dimensional varieties: cubic threefolds (Section 14). For a one-parameter family of smooth cubic threefolds in $\mathbb{P}^{4}$, the Gauss–Manin connection gives a first-order system on $H^{3}$ of rank $10$. Our formalism again produces a canonical second-order system, now on the rank-$5$ Hodge bundle $E=H^{2,1}$, showing that the same Schwarzian viewpoint extends beyond curves to a higher-dimensional variation of Hodge structure.
@@
    Mass–spring systems (Section 15). Mass–spring systems naturally give rise to second-order ordinary differential equations with matrix-valued coefficients. They provide a concrete mechanical model for the same formalism: the coefficients are most naturally viewed as geometric data on a time curve, transforming coherently under changes of local time coordinate.
@@
Part IV (Sections 12–14) treats periods of Riemann surfaces and cubic threefolds from a Schwarzian perspective.
Part V (Section 15) discusses mass–spring systems in mechanics and their relation to the Schwarzian.
@@ 2. Second order scalar ODE
The following classical computation is standard (see, for example, [12, § IV.2]). For a classical discussion of the projective viewpoint on second-order equations, see also Klein’s Lectures on the Icosahedron [18, Chapter III, §7].
@@ 4. Classical differentials and connections
Recall that $K_{X}$ denotes the canonical line bundle of holomorphic $1$-forms on $X$. Fix an integer $m$. A meromorphic connection on $K_{X}^{m}$ may be encoded locally by meromorphic functions $a_{i}\in\mathcal{M}(U_{i})$ satisfying on overlaps
@@ 5. Quantum differentials and connections
Here $I$ denotes the identity operator on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j},V)$, scalar functions act by pointwise multiplication, and the transported operator is
@@
on the open subset of $U_{i}$ where $\Psi_{i}(f)$ is pointwise invertible. Then $A=\{A_{i}\}$ is a quantum connection of eccentricity $1$ with values in $\operatorname{End}(V)$.
@@
one obtains a scalar quantum connection of eccentricity $1$.
@@ 6. Curvature
Let $A$ be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. For $m\in\mathbb{Z}$ and a quantum $m$-differential $\Psi$ with values in $V$, define the quantum covariant derivative $\nabla_{A}\Psi$ locally by
@@
If $A$ has eccentricity $e\neq 0$ and $\Psi$ is a quantum $m$-differential, then $\nabla_{A}\Psi$ is a quantum $(m+1)$-differential.
@@
Even if $A$ and $\Psi$ are classical in the sense of the previous example, the operator $\nabla_{A}\Psi$ need not be classical: the derivative term $\Psi_{i}(f)^{\prime}$ typically produces an $f^{\prime}$-term. In the classical setting one instead defines the coefficient-level covariant derivative by
@@
We briefly recall the classical Cartan structural equation for a connection with values in a Lie algebra. Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A $\mathfrak{g}$-valued connection $1$-form on a manifold $M$ is a $\mathfrak{g}$-valued differential $1$-form $A\in\Omega^{1}(M,\mathfrak{g})$. Its curvature is the $\mathfrak{g}$-valued $2$-form
@@
Let $A$ be a quantum connection on $X$ with values in $\operatorname{End}(V)$ and eccentricity $e\neq 0$. The quantum curvature of $A$ is the family of local operators
@@
Let $A$ be a quantum connection of eccentricity $e\neq 0$ with values in $\operatorname{End}(V)$. On an overlap $U_{i}\cap U_{j}$ with change of coordinate $\lambda=\lambda_{ij}$, one has
@@
where $S(\lambda)$ is the Schwarzian derivative of $\lambda$, and $I$ is the identity endomorphism.
@@
Consider the scalar quantum $1$-differential $\Psi(f)=f^{\prime}$. On the open set where $f^{\prime}\neq 0$, the Maurer–Cartan construction yields the scalar quantum connection $A(f)=f^{\prime\prime}/f^{\prime}$. Its quantum curvature is
@@
Let $A$ be a quantum connection with values in $\operatorname{End}(V)$ and eccentricity $e$. For $g\in G$ define
@@
We recall that the appearance of an affine “connection term” in the transformation law is classical; see Klein [18, Chapter III, §7] for the corresponding discussion in the genus one setting.
@@
If $A$ is a quantum connection of eccentricity $e$, then $g\bullet A$ is again a quantum connection of eccentricity $e$. Moreover, $(g,A)\mapsto g\bullet A$ defines an action of $G$.
@@
Here $I$ is the identity operator, the matrices $g_{ij}$ act by pointwise multiplication, and primes denote differentiation with respect to the local coordinate on $U_{j}$.
@@
The anomaly term $eS(\lambda)I$ is scalar (proportional to $I$). Since elements of $G$ act by conjugation and scalar endomorphisms commute with all elements of $\operatorname{End}(V)$, the relation $F_{A_{2}}=gF_{A_{1}}g^{-1}$ is independent of the choice of local coordinates. Thus the definition is well defined.
@@ 7. Second order quantum ODE
whose coefficients depend on the chosen coordinate. In the quantum setting we take the solution to be a quantum 0-differential $\Psi$, and we allow the coefficients to be quantum tensors (connections and $2$-differentials) so that the notion becomes intrinsic on $X$.
@@
as operators on $\mathcal{M}_{\mathrm{gen}}(U_{i}\cap U_{j})$. In particular, $q$ is a quantum $2$-differential, while $p$ transforms as a quantum connection of eccentricity $1$.
@@
By Lemma 7.1, the coefficient $p$ transforms as a quantum connection of eccentricity $1$. Hence we may view $p$ itself as a connection and form the corresponding covariant derivative $\nabla_{p}$ acting on quantum differentials.
@@
be the scalar quantum connection (of eccentricity $1$) on $U_{i}$. Let
@@
be the quantum $2$-differential given by squaring the first derivative. Fix a constant $\omega\in\mathbb{C}$, and set
@@
where $\Psi$ is a quantum 0-differential and $q$ is a quantum $2$-differential. Equivalently, since for $e=\tfrac{1}{2}$ and $m=0$ one has
@@ 7.1. Main Lemma
We are now ready to state our main lemma. Suppose we have an invertible solution $g$ of this equation with values in $\operatorname{End}(V)$, then the corresponding gauge transform of $A$ has curvature $F_{A}-q$ up to conjugation.
@@
Let $A$ be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. Let $g$ be an $\operatorname{End}(V)$-valued quantum 0-differential such that for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$ the meromorphic function $g_{i}(f)\in\mathcal{M}(U_{i},\operatorname{End}(V))$ satisfies $\det(g_{i}(f))\not\equiv 0$ on $U_{i}$. Assume that for every $i$ and every $f\in\mathcal{M}(U_{i})$ one has, on the open subset where $g_{i}(f)$ is pointwise invertible,
@@
Let $A$ be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. For the quantum second-order equation
@@
while $q$ transforms tensorially as a quantum $2$-differential. Hence
@@
In particular, $\mathcal{S}_{A,q}$ has the same anomaly under coordinate change as the curvature (up to the normalizing factor $2$).
@@ 8. The gauge and $\star$ actions
Let $A$ be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. The second-order equation (7.0) depends on the choice of the pair $(A,q)$, while its quantum Schwarzian $\mathcal{S}_{A,q}=2(F_{A}-q)$ is a more intrinsic quantity. We now describe a natural action of quantum $1$-differentials on such pairs which preserves $F_{A}-q$ (and hence $\mathcal{S}_{A,q}$).
Compare the scalar Miura transformation for affine and projective connections in Frenkel [8], §8.5, pp. 95–96. To the best of our knowledge, however, the following matrix-valued affine $\star$-action on pairs $(A,q)$ is new.
@@
Let $u$ be an $\operatorname{End}(V)$-valued quantum $1$-differential on $X$. We define
@@
The operation $\star$ defines an action of the additive group of $\operatorname{End}(V)$-valued quantum $1$-differentials on the set of pairs $(A,q)$ with $e$. Equivalently, for quantum $1$-differentials $u$ and $v$ one has
@@
This is a direct computation using bilinearity of the products and linearity of the derivative. Writing $(\widetilde{A},\widetilde{q}):=u\star(A,q)$, we have $\widetilde{A}=A+u$ and $\widetilde{q}=q-uA-Au-u^{2}+u^{\prime}$. Hence
[…]
@@ 8.1. Main Theorem
Let $g\in G$ be a gauge transformation (Definition 6.7). Let $A$ be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. Define
@@
Then $A^{g}$ is again a quantum connection of eccentricity $e=\tfrac{1}{2}$, and $q^{g}$ is an $\operatorname{End}(V)$-valued quantum $2$-differential. Moreover, if $\Psi$ is a (quantum) 0-differential solving
@@
and subtracting $q^{g}$ yields $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$. Multiplying by $2$ gives the stated transformation of $\mathcal{S}_{A,q}$. Conjugation invariance of characteristic polynomial coefficients and traces is immediate. ∎
@@
It is important that the transformation of $q$ in Theorem 8.4 is not given by tensorial conjugation alone. In general the curvature $F_{A}=A^{\prime}-A^{2}$ is not gauge covariant under the action $A\mapsto g\bullet A=gAg^{-1}+g^{\prime}g^{-1}$ when $A$ is $\operatorname{End}(V)$-valued and noncommutative. The additional terms $2g^{-1}Ag^{\prime}-g^{-1}g^{\prime\prime}$ in $q^{g}$ are exactly what is needed so that the Schwarzian combination $F_{A}-q$ transforms by conjugation, and hence $\mathcal{S}_{A,q}=2(F_{A}-q)$ is the natural gauge-theoretic invariant of the second order equation.
@@
Let $A$ be an $\operatorname{End}(V)$-valued quantum connection of eccentricity $e=\tfrac{1}{2}$ and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. Let $u$ be an $\operatorname{End}(V)$-valued quantum $1$-differential and let $g\in G$ be a gauge transformation. Set $u^{g}:=g^{-1}ug$, and let $(A^{g},q^{g})$ denote the gauge transform of $(A,q)$ as in Theorem 8.4. Then the $\star$ and gauge actions are compatible in the sense that
$(u\star(A,q))^{g}\;=\;u^{g}\star(A^{g},q^{g}),$
where $(\cdot)^{g}$ denotes the gauge transformation of pairs given by Theorem 8.4. Equivalently, gauge transformations send $\star$-orbits of pairs $(A,q)$ to $\star$-orbits.
@@
Write $(A^,q^)u\star$$(\widehat{A},\widehat{q})$, so $\widehat{A}=A+u$ and $\widehat{q}=q-uA-Au-u^{2}+u^{\prime}$. The first components agree:
$\widehat{A}^{g}=g^{-1}(A+u)g-g^{-1}g^{\prime}=(g^{-1}Ag-g^{-1}g^{\prime})+g^{-1}ug=A^{g}+u^{g}.$
For the second component, it is convenient to use the gauge-covariant quantity $F_{A}-q$. By Lemma 8.3 one has $F_{\widehat{A}}-\widehat{q}=F_{A}-q$, hence Theorem 8.4 gives
$F_{\widehat{A}^{g}}-\widehat{q}^{g}=g^{-1}(F_{A}-q)g.$
On the other hand, Theorem 8.4 implies $F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g$. Apply Lemma 8.3 to the pair $(A^{g},q^{g})$ and the $1$-differential $u^{g}$ to write $u^{g}\star(A^{g},q^{g})=(A^{g}+u^{g},\widehat{q}_{g})$ with
$F_{A^{g}+u^{g}}-\widehat{q}_{g}=F_{A^{g}}-q^{g}=g^{-1}(F_{A}-q)g.$
Since $\widehat{A}^{g}=A^{g}+u^{g}$, comparing the two displayed identities shows $\widehat{q}^{g}=\widehat{q}_{g}$, which proves $(u\star(A,q))^{g}=u^{g}\star(A^{g},q^{g})$. ∎
@@
The action $\star$ is transitive on the choice of the connection component: if $A_{1}$ and $A_{2}$ are two eccentricity $e=\tfrac{1}{2}$ quantum connections, then $u:=A_{2}-A_{1}$ is a quantum $1$-differential and $A_{2}=A_{1}+u$. Thus $\star$ parametrizes different presentations $(A,q)$ of the same quantum Schwarzian. In the classical scalar case this recovers the familiar fact that a projective connection can be represented by an affine connection modulo addition of a $1$-form.
@@
Let $A$ be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. Suppose $g$ is an $\operatorname{End}(V)$-valued quantum 0-differential such that $\det(g_{i}(f))\not\equiv 0$ for every chart $U_{i}$ and every $f\in\mathcal{M}(U_{i})$, and assume that $g$ satisfies the second-order equation
@@
Let $A$ be a quantum connection of eccentricity $e=\tfrac{1}{2}$ with values in $\operatorname{End}(V)$, and let $q$ be an $\operatorname{End}(V)$-valued quantum $2$-differential. Let $g_{1}$ and $g_{2}$ be two invertible (classical) $\mathrm{GL}(V)$-valued solutions of